OpenStudy (anonymous):

f(x) =|x+1|+|x-1|-3 can someone help! Is the domain. All realnumbers? Nd the range?

OpenStudy (anonymous):

the domain is all real numbers because you can take the absolute value of anything

OpenStudy (anonymous):

range is a different story if \(-1<x<1\) then you have \(|x+1|=x+1\) and \(|x-1|=1-x\) and so on this intevral you are looking at \(x+1-x+1-3=-1\) i.e. the function is never less than \(-1\)

OpenStudy (anonymous):

The range?

OpenStudy (anonymous):

Vvertex?

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.